if the temperature is -55'c what is the corresponding temperature on the kelvin scale?

a) 328 k b) 55k c)218 k d) 273k e)225k
Answer #1

As relation between degree Celsius and kelvKe scale

Kelvin = 273 + degree Celsius .......(1)

From (1)

K = 273 + (-55)

= 218 k .....Ans

Option -(c)
Answer #2
K=°C + 273.15 K=-55°C +273.15 K=218.15
source: Listening to prof haha
answered by: Lableh